# Safe Navigation in Uncertain Crowded Environments Using Risk Adaptive CVaR Barrier Functions

### v4

@@ V-C Risk Adaptive CVaR-BF Optimization
Building on the CVaR-BF constraint in (4) and leveraging the adaptive risk level $\beta_{k}$ defined in (28), we formulate an optimization problem that ensures probabilistic safety while minimally deviating from a nominal control input. Consider the robot operates in an environment with $N$ dynamic, uncertain obstacles. The objective is to design a controller that drives a mobile robot toward its target while avoiding dynamic obstacles, thereby achieving a “reach and avoid” task with minimal collision probability.
@@
To approximate the CVaR term in the safety constraint, for each obstacle, we consider a finite set of possible realizations of the next obstacle state, $\mathcal{X}^{o}_{k+1}=\{\mathbf{x}^{o,(1)}_{k+1},\mathbf{x}^{o,(2)}_{k+1},\dots,\mathbf{x}^{o,(L)}_{k+1}\}$, each associated with a probability $p(\mathbf{x}^{o,(j)}_{k+1})$. These samples may be drawn from any predictive distribution or model that captures the uncertainty in the obstacle’s next state. For each realization, we compute the corresponding value of the random variable $h_{k+1}$, which represents the safety measure at time $k+1$. These samples are then used to approximate the expectation in the CVaR formulation in (8). Using the CVaR definition from (8) and its reformulation in (13), the following optimization problem is solved at each time $k$:
@@ VI-A2 Agent Settings
We assess our method in a widely used crowd navigation simulator within a 12m $\times$ 12m space [26] (see Fig. 3), where each obstacle follows the Social Force Model (SFM) [27] with uncooperative behavior, meaning obstacles avoid collisions only with one another, not with the robot. Their maximum speeds along each axis are chosen uniformly from $\{0.3,0.6,0.9,1.2\}$m/s, and their radii are selected from $\{0.3,0.4,0.5\}$m. The robot is modeled as a double integrator, with dynamics given by
@@ VI-A3 Performance Metrics
Let $m_{t},m_{s},m_{f},$ and $m_{c}$ denote the total, successful, and feasible test cases, and the number of collisions, respectively, and let $\mathcal{M}$ be the set of successful cases. For each $i\in\mathcal{M}$, $L_{i}$ and $T_{i}$ denote the trajectory length and execution time, respectively. We evaluate the following performance metrics: 1) Success Rate (SR): $\frac{m_{s}}{m_{t}}$, 2) Feasibility Rate (FR): $\frac{m_{f}}{m_{t}}$, 3) Collision Rate (CR): $\frac{m_{c}}{m_{t}}$ 4) Average Trajectory Length (ATL): $\frac{1}{m_{s}}\sum\nolimits_{i\in\mathcal{M}}L_{i}$, 5) Average Trajectory Time (ATT): $\frac{1}{m_{s}}\sum\nolimits_{i\in\mathcal{M}}T_{i}$. A test case is considered feasible if, for every time step $k$ along the planned trajectory, the feasible set $\mathcal{F}_{k}$ (17) of the optimization problem in (LABEL:eq:opt3) is nonempty, i.e., $\forall k,\;\mathcal{F}_{k}\neq\emptyset$. Furthermore, a test case is deemed successful if it is feasible and the robot reaches the goal point, denoted by $\mathbf{x}_{\text{goal}}$, i.e., $\bigl{(}\forall k,\quad\mathcal{F}_{k}\neq\emptyset\bigr{)}\,\wedge\,\bigl{(}\exists\,K:\,\|\mathbf{x}_{k}-\mathbf{x}_{\text{goal}}\|<\varepsilon,\,\forall k\geq K\bigr{)}$, where $\varepsilon$ is a small positive threshold.
@@ VI-B Benchmark Comparisons
    CVaR-BF Methods: distance-based CVaR-BF with fixed risk levels ($\beta=0.01$ or $0.99$[5] (CVaR(Dist)).
@@
    CBF Methods: collision cone-based approaches [9], including a robust control for worst-case scenarios (RCBF(Cone)) and a standard CBF-based control without explicit uncertainty handling (CBF(Cone)).
@@
Note that both ORCA and the RL methods assume a holonomic model, resulting in a simpler problem setup compared to the other methods. For all experiment, we average on 120 different configurations for the obstacles to get the results.
@@
Table I presents the results for the scenario with $5$ obstacles and a noise level $\sigma$ = 0.05. As we can see, our proposed method achieves the highest success rate by achieving the highest feasibility rate while keeping a low collision rate, demonstrating our adaptive CVaR-BF helps improve optimization feasibility and safe decision. In contrast, the CVaR(Dist) methods show limitations: with $\beta=0.99$, the absence of a dynamic zone results in a higher collision rate, whereas with $\beta=0.01$, the lack of adaptive risk adjustment significantly reduces the feasibility rate. The cone-based methods, RCBF(Cone) and CBF(Cone), while yielding a lower collision rate, is overly conservative and also suffer from poor feasibility. Although CrowdNav++ achieves the second highest success rate, it lacks a probabilistic safety guarantee, leading to a high collision rate.
We provide more analysis of the statistical SR in Fig. 4 with respect to different number of obstacles and noise levels. Fig. 4a shows that in deterministic environment ($\sigma=0.0$), CrowdNav++ achieves the highest success rate since it is trained in the same deterministic environment. CrowdNav lacks obstacle trajectory prediction, thus exhibiting a lower success rate. Note that all CVaR-BF based methods reduce to standard CBF methods when $\sigma=0.0$, leading to ineffective risk adjustments, therefore does not show superior performance. However, since our approach incorporates a dynamic zone, it still outperforms cone-based and distance-based barrier functions. As shown in Fig. 4b, introducing noise reduces the SR of RL methods regardless of the number of obstacles. Although CrowdNav++ considers randomized obstacle behaviors (e.g., goal change and size variation), it still experiences a significant performance drop under uncertainty. More conservative methods, e.g., RCBF(Cone) and CVaR(Dist) with $\beta=0.01$ also suffer significant degradation due to feasibility issues, resulting in the worst performance. An interesting finding is that our method performs even better in uncertain environments (Fig. 4b) than in deterministic ones (Fig. 4a). This is because in deterministic environments, our method degrades to CBF, thus the entire risk adaptation mechanism becomes void, making it less flexible. Fig. 4b and Fig. 4c further indicate that as noise level increases, the performance of all methods deteriorates. However, our proposed approach consistently achieves the highest success rate, demonstrating its robustness against uncertainty.
@@ VI-C Ablation Studies
Method Setting SR CR FR ATL ATT
Proposed 0.917 0.058 1.000 13.036 13.092
w/o dynamic zone Dist-based 0.859 0.133 1.000 13.198 13.914
Cone-based 0.548 0.067 0.615 13.529 11.793
w/o risk adaptation Fix $\beta$=0.01 0.681 0.007 0.689 13.395 13.000
Fix $\beta$=0.99 0.741 0.052 0.793 13.253 12.566

### v5

@@ V-C Risk Adaptive CVaR-BF Optimization
Building on the CVaR-BF constraint in (4) and leveraging the adaptive risk level $\beta_{k}$ defined in (27), we formulate an optimization problem that ensures probabilistic safety while minimally deviating from a nominal control input. Consider the robot operates in an environment with $N$ dynamic, uncertain obstacles. The objective is to design a controller that drives a mobile robot toward its target while avoiding dynamic obstacles, thereby achieving a “reach and avoid" task with minimal collision probability.
@@
To approximate the CVaR term in the safety constraint, for each obstacle, we consider a finite set of possible realizations of the next obstacle state, $\mathcal{X}^{o}_{k+1}=(\mathbf{x}^{o,(1)}_{k+1},\mathbf{x}^{o,(2)}_{k+1},\dots,\mathbf{x}^{o,(L)}_{k+1})$, where $L$ denotes the total number of samples. Each realization is associated with a probability $p(\mathbf{x}^{o,(j)}_{k+1})$. These samples may be drawn from any predictive distribution or model that captures the uncertainty in the obstacle’s next state. For each realization, we compute the corresponding value of the random variable $h_{k+1}$, which represents the safety measure at time $k+1$. These samples are then used to approximate the expectation in the CVaR formulation in (8). Using the CVaR definition from (8) and its reformulation in (13), the following optimization problem is solved at each time $k$:
@@ VI-A3 Performance Metrics
Let $m_{t},m_{s},m_{f},$ and $m_{c}$ denote the total, successful, and feasible test cases, and the number of collisions, respectively, and let $\mathcal{M}$ be the set of successful cases. A test case is considered feasible if, for every time step $k$ along the planned trajectory, the feasible set $\mathcal{F}_{k}$ (17) of the optimization problem in (LABEL:eq:opt3) is nonempty, i.e., $\forall k,\;\mathcal{F}_{k}\neq\emptyset$. Furthermore, a test case is deemed successful if it is feasible and the robot reaches the goal point, denoted by $\mathbf{x}_{\text{goal}}$, i.e., $\bigl{(}\forall k,\quad\mathcal{F}_{k}\neq\emptyset\bigr{)}\,\wedge\,\bigl{(}\exists\,K:\,\|\mathbf{x}_{k}-\mathbf{x}_{\text{goal}}\|<\varepsilon,\,\forall k\geq K\bigr{)}$, where $\varepsilon$ is a small positive threshold. For each $i\in\mathcal{M}$, $L_{i}$ and $T_{i}$ denote the trajectory length and execution time, respectively. We evaluate the following performance metrics: 1) Success Rate (SR): $\frac{m_{s}}{m_{t}}$, 2) Feasibility Rate (FR): $\frac{m_{f}}{m_{t}}$, 3) Collision Rate (CR): $\frac{m_{c}}{m_{t}}$ 4) Average Trajectory Length (ATL): $\frac{1}{m_{s}}\sum\nolimits_{i\in\mathcal{M}}L_{i}$, 5) Average Trajectory Time (ATT): $\frac{1}{m_{s}}\sum\nolimits_{i\in\mathcal{M}}T_{i}$.
@@ VI-B Benchmark Comparisons
    CVaR-BF Methods: distance-based CVaR-BF with fixed risk levels ($\beta=0.01$ or 0.990.990.99[6] (CVaR(Dist)).
@@
    CBF Methods: collision cone-based CBF, as defined in (LABEL:eq:cone[11], including a robust control for worst-case scenarios (RCBF(Cone)) and a standard CBF-based control without explicit uncertainty handling (CBF(Cone)).
@@
Note that both ORCA and the RL methods assume a holonomic model, resulting in a simpler problem setup compared to the other methods. For all experiments, we average on 120 different configurations for the obstacles to get the results.
@@
Table I presents the results for the scenario with 555 obstacles and a noise level $\sigma$ = 0.05. As we can see, our proposed method achieves the highest success rate by achieving the highest feasibility rate while keeping a low collision rate, demonstrating our adaptive CVaR-BF helps improve optimization feasibility and safe decision. In contrast, the CVaR(Dist) methods show limitations: with $\beta=0.99$, the absence of a dynamic zone results in a higher collision rate, whereas with $\beta=0.01$, the lack of adaptive risk adjustment significantly reduces the feasibility rate. The cone-based methods, RCBF(Cone) and CBF(Cone), while yielding a lower collision rate, are overly conservative and also suffer from poor feasibility. Although CrowdNav++ achieves the second highest success rate, it lacks a probabilistic safety guarantees, leading to a high collision rate.
We provide a more detailed analysis of the statistical SR in Fig. 4 with respect to different number of obstacles and noise levels. Fig. 4a shows that in a deterministic environment ($\sigma=0.0$), CrowdNav++ achieves the highest success rate since it is trained in the same deterministic environment. CrowdNav lacks obstacle trajectory prediction, thus exhibiting a lower success rate. Note that all CVaR-BF based methods reduce to standard CBF methods when $\sigma=0.0$, leading to ineffective risk adjustments, therefore do not show superior performance. However, since our approach incorporates a dynamic zone, it still outperforms the cone-based and distance-based barrier functions. As shown in Fig. 4b, introducing noise reduces the SR of RL methods regardless of the number of obstacles. Although CrowdNav++ considers randomized obstacle behaviors (e.g., goal change and size variation), it still experiences a significant performance drop under uncertainty. More conservative methods, e.g., RCBF(Cone) and CVaR(Dist) with $\beta=0.01$ also suffer significant degradation due to feasibility issues, resulting in the worst performance. An interesting finding is that our method performs even better in uncertain environments (Fig. 4b) than in deterministic ones (Fig. 4a). This is because in uncertain environments, the risk adaptation mechanism becomes effective and provides additional flexibility. Fig. 4b and Fig. 4c further indicate that as noise level increases, the performance of all methods deteriorates. However, our proposed approach consistently achieves the highest success rate, demonstrating its robustness against uncertainty.
@@ VI-C2 Effectiveness Validation of Risk Adaptation
Method Setting SR CR FR ATL ATT
Proposed 0.917 0.058 1.000 13.036 13.092
w/o
dynamic zone
Dist-based 0.859 0.133 1.000 13.198 13.914
Cone-based 0.548 0.067 0.615 13.529 11.793
w/o
risk adaptation
Fix $\beta$=0.01 0.681 0.007 0.689 13.395 13.000
Fix $\beta$=0.99 0.741 0.052 0.793 13.253 12.566